### v1

@@ Abstract
Figure 2: Common image classification tasks are invariant to flopping (horizontal mirroring). In our implementation, this invariance is enforced through equivariant network layers, reducing the required floating-point operations (FLOPs) by half.
@@ 1 Introduction
One of the main drivers of progress in deep learning is the scaling of compute. This idea is perhaps best summarized in Sutton’s Bitter Lesson Sutton (2019). As Sutton states:
@@
It follows from the Bitter Lesson that a guiding principle for designing deep learning models should be to find simple models that scale well. In computer vision, this has proven fruitful and led to models such as the Vision Transformer (ViT) Dosovitskiy et al. (2021). However, a straightforward critique of Sutton’s argument is that leveraging domain knowledge can yield algorithms that better utilize computation. This paper explores simple image symmetries that enable more efficient computations. In a nutshell, our case is that equivariant neural networks are simple models that scale well.
Many vision problems are invariant (or more generally equivariant) to horizontal mirroring, also known as flopping [^1] See Wikipedia article or Oxford definition., and partially invariant to other geometric transformations such as rotations. For instance, both images in Figure 2 show a butterfly of the same species. Enforcing geometric invariance in a neural network architecture can be done by hard-coding the symmetry in each network layer, by enforcing specific weight sharing Wood & Shawe-Taylor (1996); Cohen & Welling (2016). It has been shown in the literature that such architectural symmetry constraints typically improve the parameter efficiency of neural networks Cohen & Welling (2016); Bekkers et al. (2018); Weiler & Cesa (2019). However, this comes at the cost of requiring more compute per parameter. For instance, Klee et al. (2023) found that rotation invariant ResNets typically outperformed ordinary ResNets on ImageNet-1K Deng et al. (2009); Russakovsky et al. (2015); Recht et al. (2019) in the equal-parameter setting, but that the training time of the invariant ResNets far exceeded that of the ordinary ResNets. This is because the symmetry-induced weight sharing means that each trainable parameter is used in more computations.
@@
Section 2 reviews related work on equivariant networks and Section 3 provides an exposition of flopping-equivariant networks for image classification, while Section A lays out the underlying mathematical background in detail. In Section 4, we design flopping-equivariant versions of popular modern vision architectures. Experiments in Section 6 on ImageNet-1K Deng et al. (2009); Russakovsky et al. (2015); Recht et al. (2019) demonstrate that as network size scales up, flopping-equivariant models achieve comparable or improved classification accuracy for ResMLPs  Touvron et al. (2023), ConvNeXts Liu et al. (2022) and ViTs Dosovitskiy et al. (2021), while requiring only half the FLOPs.
@@ 2 Related work
Our work is part of the broad field of geometric deep learning Bronstein et al. (2021). In this section, we aim to discuss the most relevant work to the present paper, but we reference many other relevant works throughout.
@@ 2.1 Equivariant networks for image input
Most work on equivariant networks on image input concerns convolutional networks (ConvNets) Fukushima (1975); LeCun et al. (1989). ConvNets are themselves equivariant to (cyclic) image translations. The ConvNets in this paper are built on ConvNeXt Liu et al. (2022), which is a modern variant with state-of-the-art classification performance.
Making ConvNets equivariant to rotations and reflections was done by Cohen & Welling (2016) and Dieleman et al. (2016), with many follow-up works. The unifying framework of steerable $E(2)$-equivariant ConvNets Cohen & Welling (2017); Weiler & Cesa (2019), subsumes much of the prior and subsequent work and proposes decomposing the feature spaces into irreps (irreducible representations, see Section A) of the symmetry group. Our ConvNeXt-variants are special cases of this general framework too.
Apart from ConvNets, there are works proposing attention- and transformer-based equivariant vision architectures Romero et al. (2020); Xu et al. (2023). Most similar to our ViT-based networks are the $\mathrm{SO}(2)$-steerable transformers by Kundu & Kondor (2024). The main differences are the group considered, the fact that they use complex-valued features, the nonlinearities used and the type of positional embeddings. However, the main idea of parametrizing the features in terms of irreps (following steerable ConvNets) and modifying the ViT architecture accordingly is the same in our ViTs and the one by Kundu & Kondor (2024).
@@ 2.2 Equivariant networks for other input than images
Brehmer et al. (2024) found that the equivariant transformer GATr Brehmer et al. (2023) outperformed non-equivariant transformers on a task of 3D rigid-body interactions, both in terms of parameter-efficiency and learning-efficiency, i.e., the number of FLOPs required during training. Unlike our models, the number of FLOPs-per-parameter remained higher than that of the non-equivariant transformer.
Bekkers et al. (2024) proposed a fast equivariant group convolutional network on point cloud input by using separable group convolutions on position orientation space $\mathbb{R}^{3}\times S^{2}$ and parallelizing the message passing step. Their networks use ConvNeXt-style blocks, as do some of our networks.
@@ 3.1 Linear layers
Linear layers of the form (2) are called equivariant. In general, any layer that preserves the transformation properties of the input is called equivariant. For instance, we can add a bias parameter to the equivariant linear layer if we enforce the bias to be zero for the $(-1)$-equivariant features. We give a more formal definition of equivariance in Section A, cf. (11). The stacking of equivariant layers to build a symmetry-respecting network is sometimes called the geometric deep learning blueprint Bronstein et al. (2021), which goes back at least to the 90’s with the works by Wood & Shawe-Taylor (1996). Using flopping invariant and $(-1)$-equivariant features is a special case of the steerable features proposed by Cohen & Welling (2017). In the networks by Cohen & Welling (2017), the described computational boost is not realized, as they implement their equivariant convolutions in terms of ordinary convolutions. We will discuss this in more depth in Section 4.4.
@@ 3.2 Patch embedding layer
Obtaining invariant and $(-1)$-equivariant features from an image is easy. One straightforward manner is to convolve the image with symmetric and antisymmetric filters as illustrated in Figure 3. Such convolutions can be incorporated as the first layer in convolutional neural networks and vision transformers alike (the so-called patch embedding layer, or “PatchEmbed”).
Figure 3: The patch embedding (PatchEmbed) layer is common in modern networks, proposed with the ViT-model Dosovitskiy et al. (2021). PatchEmbed is a convolutional layer with stride equal to the kernel size of the convolution filters. In our equivariant architectures, we enforce half of the filters to be symmetric and half to be antisymmetric. When the input image is flopped, the output feature map of a symmetric filter is flopped as well. The output feature map of an antisymmetric filter is flopped and changes sign. Half of the features output from our PatchEmbed layer are flopping invariant, while half are $(-1)$-equivariant, enabling efficient equivariant processing in subsequent linear layers, as detailed in Section 3.1.
@@ 3.3 Non-linear layers
Non-linear layers of the neural network also need to be equivariant, i.e. respect the transformation rules of the features. For instance, common layers in modern networks include layer normalization Ba et al. (2016), activation functions such as GELU Hendrycks & Gimpel (2016), as well as multi-head self-attention Bahdanau et al. (2015); Vaswani et al. (2017). Layer normalization is equivariant without modification since the norm of neither $x_{1}$ nor $x_{-1}$ changes as the input is flopped. For pointwise activation functions $\sigma$, we can create equivariant versions by computing the output $y_{1},y_{-1}$ through
@@ 3.5 Limitations of equivariant networks
Finally, we need to discuss the limitations put on the model by enforcing each layer to be equivariant. Prior work has shown experimentally, and theoretically in some special cases, that ordinary networks learn to be approximately equivariant by training on symmetric data Lenc & Vedaldi (2015); Olah et al. (2020); Gruver et al. (2023); Bruintjes et al. (2023); Bökman & Kahl (2023); Marchetti et al. (2024). What is meant by this is that for a given output feature space of a particular layer in the network, there exists a “steering matrix” $A$, such that when the input is flopped, the output feature is (approximately) transformed by $A$. Since flopping an image twice returns the original image, we must have that $A^{2}=I$, which means that $A$ can be diagonalized as $A=QDQ^{-1}$ with $D$ diagonal containing only entries of ±$\pm 1$. (The generalization of this diagonalization is called the isotypical decomposition and is possible for more general symmetry groups as well, due to Maschke’s theorem (Theorem A.1).) In other words, by changing the basis that we parametrize the network features in by $Q$, we get features that are invariant and $(-1)$-equivariant as before. Therefore, restricting the network to be equivariant is not really limiting what features are learnt since the equivariance is learned even without this restriction.
@@ 4 Modern neural networks for image data
Since the development of Highway nets Srivastava et al. (2015) and ResNets He et al. (2016a), modern neural networks for image data consist of stacking residual blocks
@@
We cover three architectures based on (5), and outline the minimal changes to make them flopping-equivariant. We start with the simplest architecture, ResMLP Touvron et al. (2023), working our way through ViT Dosovitskiy et al. (2021) and finally ConvNeXt  Liu et al. (2022).
@@ 4.1 ResMLP
One unique aspect of ResMLP is that it doesn’t normalize features, but instead uses an affine layer (over the channel dimension)
@@
Figure 4: Our architectures follow the basic geometric deep learning blueprint Wood & Shawe-Taylor (1996); Cohen & Welling (2016); Bronstein et al. (2021) in combination with modern vision models Dosovitskiy et al. (2021); Liu et al. (2022); Touvron et al. (2023). The equivariant PatchEmbed is described in Section 3.2. The residual blocks differ per architecture and are explained in Section 4. The classifier head is a single linear layer.
@@ 4.2 ViT
Vision transformers (ViTs) use transformer encoder blocks Vaswani et al. (2017) as residual blocks. The layer closely follows the original transformer implementation and only deviates by the layer normalization being applied before the computation, as proposed by He et al. (2016b).
@@
To make ViT flopping-equivariant we further modify the positional encoding to consist of half channels that are symmetric about the middle of the image and half channels that are anti-symmetric, thus forming invariant and $(-1)$-equivariant features. Positional encodings are added to the features after the initial PatchEmbed layer. We also enforce the classification token appended to the embedded image patches in ViT to be constant zero for the $(-1)$-equivariant features. Multi-head self attention is made equivariant by using block-diagonal linear layers for the projections to queries, keys and values followed by attention as described in Section 3.4.
Following Weiler & Cesa (2019), we also implement a hybrid model333 These are called “restricted” by Weiler & Cesa (2019). where the first half of the residual blocks are equivariant and the second half are standard blocks. The intuition is that inductive bias is most useful in early network layers for learning general features in all orientations, while the last layers use more specialized processing. As this is not the focus of the paper, we include only hybrid versions of the ViTs, denoted $\mathcal{H}(\text{ViT})$, to test the limits of scaling equivariant layers.
@@ 4.4 On efficient steerable convolutions
In prior work Weiler & Cesa (2019), steerable convolutions were implemented by inputting equivariant filters in ordinary convolution software routines. This did not provide any computational benefit over non-equivariant filters, which use the same routines. Our implementations are more efficient because the ConvNeXt architecture utilizes depthwise convolutions followed by $1\times 1$-convolutions, and the $1\times 1$-convolutions are pixel-wise linear layers that can be decomposed into block-diagonals as in (2). There are previous works that use separable convolutions in equivariant networks, but they parametrize the features in the spatial domain rather than in terms of irreps and do not obtain a computational advantage over ordinary separable convolutions Lengyel & van Gemert (2021); Knigge et al. (2022).
@@
Even more efficiently, one can use so-called Winograd schemes specific for symmetric or anti-symmetric filters Winograd (1980). We do not pursue this optimization in the present paper, as it requires significant effort into writing GPU-code, but we believe that it is a promising direction for further improving the runtime of steerable ConvNets. Thus, the FLOP and throughput values reported for (ConvNeXt $\mathcal{E}(\text{ConvNeXt ({iso.})})$. ) ) in Table 1 do not include the possible further reduction obtainable from more efficient symmetric and anti-symmetric depthwise convolutions.
@@
There are a couple of different ways that equivariant networks can be compared to ordinary networks. Cohen & Welling (2016) and most later works compare an equivariant network with $M$ trainable parameters with ordinary networks with $M$ parameters. This is a fair comparison from a learning theory perspective since the parameters can then store the same amount of information. Subsequent works noted that this is unfair from a computational perspective, Weiler & Cesa (2019); Klee et al. (2023); Roos & Kroon (2024), because of the fact that the equivariant ConvNets had the same computational cost as ordinary network with many more parameters, since as mentioned they used the same convolution routines.
The first measure for computational cost that we will use in this paper is the number of FLOPs required for a forward pass of the network. This represents a bound on how efficient the network can be made. In the limit, as feature spaces grow and dense linear layers dominate the compute, the number of FLOPs accurately predicts how fast the network will run on modern GPUs. FLOPs are therefore used for analyzing the compute scaling of large networks Kaplan et al. (2020). However, the most direct measure of computational cost is the network’s throughput in terms of images per second. This is a somewhat inconsistent measure as it depends heavily on the hardware used and can sometimes vary drastically with minor implementation changes. We believe that the best way to compare networks is to report all three measures: number of parameters, FLOPs and throughput, as is usually done in the literature on non-equivariant networks. One of the main points of this paper is to illustrate that contrary to popular belief, equivariant networks can be designed to have approximately the same number of FLOPs per parameter as ordinary networks (see Figure 5).

### v2

@@ Abstract
Deep Learning, Computer Vision, Equivariance, Computational Efficiency
@@ 1 Introduction
One of the main drivers of progress in deep learning is the scaling of compute. This idea is perhaps best summarized in Sutton’s Bitter Lesson (Sutton, 2019). As Sutton states:
@@
It follows from the Bitter Lesson that a guiding principle for designing deep learning models should be to find simple models that scale well. In computer vision, this has proven fruitful and led to models such as the Vision Transformer (ViT) (Dosovitskiy et al., 2021). However, a straightforward critique of Sutton’s argument is that leveraging domain knowledge can yield algorithms that better utilize computation. This paper explores common image symmetries that enable more efficient computations. In a nutshell, our case is that equivariant neural networks can be simple models that scale well.
Figure 1: Common image classification tasks are invariant to flopping (horizontal mirroring). In our implementation, this invariance is enforced through equivariant network layers, halving the required floating-point operations (FLOPs).
Many vision problems are invariant (or more generally equivariant) to horizontal mirroring, also known as flopping [^1] See Wikipedia article or Oxford definition., and partially invariant to other geometric transformations such as rotations. For instance, both images in Figure 1 show a butterfly of the same species. Enforcing geometric invariance in a neural network architecture can be done by hard-coding the symmetry in each network layer, by enforcing specific weight sharing (Wood & Shawe-Taylor, 1996; Cohen & Welling, 2016). It has been shown in the literature that such architectural symmetry constraints typically improve the parameter efficiency of neural networks (Cohen & Welling, 2016; Bekkers et al., 2018; Weiler & Cesa, 2019). However, this comes at the cost of requiring more compute per parameter. For instance, Klee et al. (2023) found that rotation invariant ResNets typically outperformed ordinary ResNets on ImageNet-1K in the equal-parameter setting, but that the training time of the invariant ResNets far exceeded that of the ordinary ResNets. This is because the symmetry-induced weight sharing means that each trainable parameter is used in more computations.
@@
Section 2 reviews related work on equivariant networks and Section 3 provides an exposition of flopping-equivariant networks for image classification, while Section A lays out the underlying mathematical background in detail. In Section 4, we design flopping-equivariant versions of popular modern vision architectures. Experiments in Section 6 on ImageNet-1K demonstrate that as network size scales up, flopping-equivariant models achieve comparable or improved classification accuracy for ResMLPs  (Touvron et al., 2023), ConvNeXts (Liu et al., 2022) and ViTs (Dosovitskiy et al., 2021), while requiring only half the FLOPs.
@@ 2 Related Work
Our work is part of the broad field of geometric deep learning (Bronstein et al., 2021). In this section, we aim to discuss the most relevant work to the present paper, but we reference many other relevant works throughout.
@@ 2.1 Equivariant Networks for Image Input
Most work on equivariant networks on image input concerns convolutional networks (ConvNets) (Fukushima, 1975; LeCun et al., 1989). ConvNets are themselves equivariant to (cyclic) image translations. The ConvNets in this paper are built on ConvNeXt (Liu et al., 2022), which is a modern variant with state-of-the-art classification performance.
Making ConvNets equivariant to rotations and reflections was done by Cohen & Welling (2016) and Dieleman et al. (2016), with many follow-up works. The unifying framework of steerable $E(2)$-equivariant ConvNets (Cohen & Welling, 2017; Weiler & Cesa, 2019), subsumes much of the prior and subsequent work and proposes decomposing the feature spaces into irreps (irreducible representations, see Section A) of the symmetry group. Our ConvNeXt-variants are special cases of this general framework too.
Apart from ConvNets, there are works proposing attention- and transformer-based equivariant vision architectures (Romero et al., 2020; Xu et al., 2023). Most similar to our ViT-based networks are the $\mathrm{SO}(2)$-steerable transformers by Kundu & Kondor (2024). The main differences are the group considered, the fact that they use complex-valued features, the nonlinearities used and the type of positional embeddings. However, the main idea of parametrizing the features in terms of irreps (following steerable ConvNets) and modifying the ViT architecture accordingly is the same in our ViTs and the one by Kundu & Kondor (2024).
@@ 2.2 Equivariant Networks for Other Input Than Images
Equivariant networks constitute a broad research direction. One early line of research is by Wood & Shawe-Taylor (1996) and going back earlier the topic connects to steerable filters (Knutsson & Granlund, 1983; Freeman & Adelson, 1991). Recent approaches include canonicalization (Kaba et al., 2023; Mondal et al., 2023), learned equivariance (Gupta et al., 2024), structured matrices (Samudre et al., 2025) et cetera. We will use the standard approach of enforcing each layer of the network to be equivariant through constraints on the weight matrices (Wood & Shawe-Taylor, 1996; Cohen & Welling, 2016).
There has been recent interest in the scaling properties of equivariant networks. Brehmer et al. (2024) find that the equivariant transformer GATr (Brehmer et al., 2023) outperforms non-equivariant transformers on a task of 3D rigid-body interactions, both in terms of parameter-efficiency and learning-efficiency, i.e., the number of FLOPs required during training.
Bekkers et al. (2024) propose a fast equivariant group convolutional network on point cloud input by using separable group convolutions on position orientation space $\mathbb{R}^{3}\times S^{2}$ and parallelizing the message passing step. Their networks use ConvNeXt-style blocks, as do some of our networks. Follow-up work extends the message passing to use universal invariants (Bellaard et al., 2025), and experimentally analyzes the benefits of equivariance in different point cloud processing tasks (Vadgama et al., 2025).
@@ 3.1 Linear Layers
Linear layers of the form (2) are called equivariant. In general, any layer that preserves the transformation properties of the input is called equivariant. For instance, we can add a bias parameter to the equivariant linear layer if we enforce the bias to be zero for the $(-1)$-equivariant features. We give a more formal definition of equivariance in Section A, cf. (11). The stacking of equivariant layers to build a symmetry-respecting network is sometimes called the geometric deep learning blueprint (Bronstein et al., 2021), which goes back at least to the 90’s with the works by Wood & Shawe-Taylor (1996). Using flopping invariant and $(-1)$-equivariant features is a special case of the steerable features proposed by Cohen & Welling (2017). However, general convolutional layers can not be be as straightforwardly block-diagonalized as in (2), since the convolutional kernels have a spatial extent. In particular, in the networks by Cohen & Welling (2017), the described computational boost is not realized, as the equivariant convolutions are implemented in terms of ordinary convolutions. We will discuss this in more depth in Section 4.4.
@@ 3.2 Patch Embedding Layer
Obtaining invariant and $(-1)$-equivariant features from an image is easy. One straightforward manner is to convolve the image with symmetric and antisymmetric filters as illustrated in Figure 2. Such convolutions can be incorporated as the first layer in convolutional neural networks and vision transformers alike (the so-called patch embedding layer, or “PatchEmbed”) and correspond to the lifting layer in an equivariant ConvNet (Cohen & Welling, 2016).
Figure 2: Patch embedding layer. The patch embedding (PatchEmbed) layer is common in modern networks, proposed with the ViT-model (Dosovitskiy et al., 2021). PatchEmbed is a convolutional layer with stride equal to the kernel size of the convolution filters. In our equivariant architectures, we enforce half of the filters to be symmetric and half to be antisymmetric. When the input image is flopped, the output feature map of a symmetric filter is flopped as well. The output feature map of an antisymmetric filter is flopped and changes sign. Half of the features output from our PatchEmbed layer are flopping invariant, while half are $(-1)$-equivariant, enabling efficient equivariant processing in subsequent linear layers, as detailed in Section 3.1.
@@ 3.3 Non-Linear Layers
Non-linear layers of the neural network also need to be equivariant, i.e. respect the transformation rules of the features. For instance, common layers in modern networks include layer normalization (Ba et al., 2016), activation functions such as GELU (Hendrycks & Gimpel, 2016), as well as multi-head self-attention (Bahdanau et al., 2015; Vaswani et al., 2017). Layer normalization is equivariant without modification since the norm of neither $x_{1}$ nor $x_{-1}$ changes as the input is flopped. For pointwise activation functions $\sigma$, we can create equivariant versions by computing the output $y_{1},y_{-1}$ through
@@ 3.5 Limitations of Equivariant Networks
Finally, we need to discuss the limitations put on the model by enforcing each layer to be equivariant. Prior work has shown experimentally, and theoretically in some special cases, that ordinary networks learn to be approximately equivariant by training on symmetric data (Lenc & Vedaldi, 2015; Olah et al., 2020; Gruver et al., 2023; Bruintjes et al., 2023; Bökman & Kahl, 2023; Marchetti et al., 2024) [^2] Some prior work also found evidence in the other direction, i.e., that weight symmetries do not always appear in networks trained on symmetric data (Moskalev et al., 2023). Thus, it is a partially unsettled question to what extent and in which situations equivariance can be learned from data. As we argue in this section, it is however the case that even if we take the least favourable view for motivating hard-coding equivariance—saying that it can be learned from data—there is still a clear computational argument in favour of hard-coding.. What is meant by this is that for a given output feature space of a particular layer in the network, there exists a “steering matrix” $A$, such that when the input is flopped, the output feature is (approximately) transformed by $A$. Since flopping an image twice returns the original image, we must have that $A^{2}=I$, which means that $A$ can be diagonalized as $A=QDQ^{-1}$ with $D$ diagonal containing only entries of ±$\pm 1$. (The generalization of this diagonalization is called the isotypical decomposition and is possible for more general symmetry groups as well, due to Maschke’s theorem (Theorem A.1).) In other words, by changing the basis that we parametrize the network features in by $Q$, we get features that are invariant and $(-1)$-equivariant as before. Therefore, restricting the network to be equivariant is not really limiting what features are learnt if the equivariance is learned even without this restriction.
@@
3.6 Generalization to Other Groups
The argument given for computational savings in linear layers can be generalized to other groups than the flopping group. In Section A, we go through the mathematical background in detail. Importantly, whenever the representations acting on the feature spaces are completely reducible (i.e. a direct sum of irreps), linear layers can always be block-diagonalized by parameterizing them in terms of the irreps. The computational savings depend on (a) the number of irreps, (b) the dimensionality of the irreps, and (c) whether the real irreps are of real, complex or quaternion type. Computational savings are possible in general but the number of FLOPs per parameter will be higher for irreps with dimension larger than $1$.
@@ 4 Modern Neural Networks for Image Data
Since the development of Highway nets (Srivastava et al., 2015) and ResNets (He et al., 2016a), modern neural networks for image data consist of stacking residual blocks
@@
We cover three architectures based on (5), and outline the minimal changes to make them flopping-equivariant. We start with the simplest architecture, ResMLP (Touvron et al., 2023), working our way through ViT (Dosovitskiy et al., 2021) and finally ConvNeXt  (Liu et al., 2022).
@@ 4.1 ResMLP
One unique aspect of ResMLP is that it doesn’t normalize features, but instead uses an affine layer (over the channel dimension), in symbols
@@
Figure 3: Schematic. Our architectures follow the basic geometric deep learning blueprint (Wood & Shawe-Taylor, 1996; Cohen & Welling, 2016; Bronstein et al., 2021) in combination with modern vision models (Dosovitskiy et al., 2021; Liu et al., 2022; Touvron et al., 2023). The equivariant PatchEmbed is described in Section 3.2. The residual blocks differ per architecture and are explained in Section 4. The classifier head is a single linear layer.
@@ 4.2 ViT
Vision transformers (ViTs) use transformer encoder blocks (Vaswani et al., 2017) as residual blocks. The layer closely follows the original transformer implementation and only deviates by the layer normalization being applied before the computation, as proposed by He et al. (2016b).
@@
To make ViT flopping-equivariant we further modify the positional encoding to consist of half channels that are symmetric about the middle of the image and half channels that are anti-symmetric, thus forming invariant and $(-1)$-equivariant features. Positional encodings are added to the features after the initial PatchEmbed layer. We also enforce the classification token appended to the embedded image patches in ViTs to be constant zero for the $(-1)$-equivariant features. Multi-head self attention is made equivariant by using block-diagonal linear layers for the projections to queries, keys and values followed by attention as described in Section 3.4.
To experiment with varying the group representations in the intermediate layers, we implement ViT-versions that use only invariant features in half the layers. After the first half of the network, we map the $d/2$ invariant dimensions linearly to $d$ invariant dimensions and throw away the $(-1)$-equivariant features. The subsequent layers are then ordinary ViT-layers, which preserve invariance. In these last layers, we do not get any computational saving compared to the baseline since the block-diagonalization in (2) will not have any $W_{-1,-1}$ or 00-blocks. We denote these half-way invariantized networks by $\mathcal{I}(\text{ViT})$.
Following Weiler & Cesa (2019), we also implement a hybrid model444 These are called “restricted” by Weiler & Cesa (2019). where the first half of the residual blocks are equivariant and the second half are standard blocks. The intuition is that inductive bias is most useful in early network layers for learning general features in all orientations, while the last layers use more specialized processing. As this is not the focus of the paper, we include only hybrid versions of the ViTs, denoted $\mathcal{H}(\text{ViT})$, to test the limits of scaling equivariant layers.
@@ 4.4 On Efficient Steerable Convolutions
In prior work (Weiler & Cesa, 2019), steerable convolutions were implemented by inputting equivariant filters in ordinary convolution software routines. This did not provide any computational benefit over non-equivariant filters, which use the same routines. Our implementations are more efficient because the ConvNeXt architecture utilizes depthwise convolutions followed by $1\times 1$-convolutions, and the $1\times 1$-convolutions are pixel-wise linear layers that can be decomposed into block-diagonals as in (2). There are previous works that use separable convolutions in equivariant networks, but they parametrize the features in the spatial domain rather than in terms of irreps and do not obtain a computational advantage over ordinary separable convolutions (Lengyel & van Gemert, 2021; Knigge et al., 2022).
@@
Even more efficiently, one can use so-called Winograd schemes specific for symmetric or anti-symmetric filters (Winograd, 1980). We do not pursue this optimization in the present paper, as it requires significant effort into writing GPU-code, but we believe that it is a promising direction for further improving the runtime of steerable ConvNets. Thus, the FLOP and throughput values reported for (ConvNeXt $\mathcal{E}(\text{ConvNeXt ({iso.})})$. ) ) in Table 1 do not include the possible further reduction obtainable from more efficient symmetric and anti-symmetric depthwise convolutions.
@@
There are a couple of different ways that equivariant networks can be compared to ordinary networks. Cohen & Welling (2016) and most later works compare an equivariant network with $M$ trainable parameters with ordinary networks with $M$ parameters. This is a fair comparison from a learning theory perspective since the parameters can then store the same amount of information. Subsequent works noted that this is unfair from a computational perspective, (Weiler & Cesa, 2019; Klee et al., 2023; Roos & Kroon, 2024), because of the fact that the equivariant ConvNets had the same computational cost as ordinary network with many more parameters, since as mentioned they used the same convolution routines.
The first measure for computational cost that we will use in this paper is the number of FLOPs required for a forward pass of the network. This represents a bound on how efficient the network can be made. In the limit, as feature spaces grow and dense linear layers dominate the compute, the number of FLOPs accurately predicts how fast the network will run on modern GPUs. FLOPs are therefore used for analyzing the compute scaling of large networks (Kaplan et al., 2020). However, the most direct measure of computational cost is the network’s throughput in terms of images per second. This is a somewhat inconsistent measure as it depends heavily on the hardware used and can sometimes vary drastically with minor implementation changes. We believe that the best way to compare networks is to report all three measures: number of parameters, FLOPs and throughput, as is usually done in the literature on non-equivariant networks. One of the main points of this paper is to illustrate that contrary to popular belief, flopping equivariant networks can be designed to have approximately the same number of FLOPs per parameter as ordinary networks (see Figure 4).